# Large Language Model Agent for Modular Task Execution in Drug Discovery

### v1

@@ Abstract
We present a modular framework powered by large language models (LLMs) that automates and streamlines key tasks across the early-stage computational drug discovery pipeline. By combining LLM reasoning with domain-specific tools, the framework performs biomedical data retrieval, domain-specific question answering, molecular generation, property prediction, property-aware molecular refinement, and 3D protein–ligand structure generation. In a case study targeting BCL-2 in lymphocytic leukemia, the agent autonomously retrieved relevant biomolecular information—including FASTA sequences, SMILES representations, and literature—and answered mechanistic questions with improved contextual accuracy over standard LLMs. It then generated chemically diverse seed molecules and predicted 67 ADMET-related properties, which guided iterative molecular refinement. Across two refinement rounds, the number of molecules with QED $>$ 0.6 increased from 34 to 55, and those passing at least four out of five empirical drug-likeness rules rose from 29 to 52, within a pool of 194 molecules. The framework also employed Boltz-2 to generate 3D protein–ligand complexes and provide rapid binding affinity estimates for candidate compounds. These results demonstrate that the approach effectively supports molecular screening, prioritization, and structure evaluation. Its modular design enables flexible integration of evolving tools and models, providing a scalable foundation for AI-assisted therapeutic discovery.
@@ 1 Introduction
The discovery of drug-like molecules for treating specific diseases is fundamental to therapeutic innovation and the advancement of public health. However, identifying viable candidates is a highly complex and resource-intensive process. It requires navigating an enormous chemical space, integrating heterogeneous biological and chemical data sources, and conducting iterative experimental validation. Collectively, these challenges contribute to drug development timelines of 10–15 years and costs exceeding \$2 billion per approved therapy 1, 2, 3.
@@
To mitigate these challenges, computational approaches have become essential in early-stage drug development. Methods such as molecular docking 4, 5, 6, quantitative structure–activity relationship (QSAR) modeling 7, 8, 9, and molecular dynamics simulations 10, 11 have significantly improved the speed and accuracy of compound screening and optimization. More recently, machine learning (ML) techniques have expanded this toolkit by enabling predictive modeling of pharmacokinetic and pharmacodynamic properties, de novo molecular generation, and multitask learning across diverse biomedical tasks12, 13, 14, 15. Notably, AlphaFold has achieved high-accuracy protein structure prediction 16, 17, 18, while Graph Neural Networks (GNNs) and transformer-based models improve property prediction 19, 20, 21, 22. In parallel, generative models such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and diffusion models (e.g., JT-VAE, MolGAN, MolDiff) enable de novo or target-specific molecular design 23, 24, 25. Collectively, these methods increase the scalability and automation potential of modern drug discovery workflows.
@@
Despite this progress, most computational methods remain task-specific and require manual orchestration by domain experts. However, drug discovery is inherently a multi-step, interdependent process that requires seamless integration across diverse tasks 1, 2, 3. This fragmented implementation limits the scalability and efficiency of current pipelines. As the demand for faster and more cost-effective drug screening grows, there is an urgent need for unified, intelligent platforms capable of autonomously coordinating these tasks while supporting expert decision-making.
@@
Large Language Models (LLMs) offer a compelling solution to this need. Trained on massive corpora of natural language data, LLMs exhibit strong reasoning capabilities and domain-agnostic knowledge. When augmented with external tools, such as domain-specific plugins, APIs, and software libraries, LLM-based agents can overcome the limitations of general-purpose language models and act as interpretable, flexible controllers of scientific workflows 26. LLM agents have recently been applied to automate diverse aspects of scientific discovery, including experimental design, material synthesis planning, and data analysis 27, 28, 29, 30, 31. For instance, Boiko et al. introduced Coscientist, an LLM agent capable of autonomously planning and executing chemistry experiments, significantly enhancing lab productivity while reducing human intervention 32. Similarly, the dZiner framework leverages LLM agents for molecular design through iterative reasoning and structure-based optimization33.
@@ 2.1 Task Modules
Data Extraction. AgentD is capable of retrieving biomedical information from both structured and unstructured sources. Given a query - such as identifying known drugs associated with a specific protein–disease context - the agent retrieves the protein’s FASTA sequence from the UniProt database 34, searches for relevant drug names via the web, and extracts corresponding SMILES representations from the ChEMBL database 35. Additionally, AgentD autonomously constructs keyword-based queries to download relevant open-access scientific literature from Semantic Scholar, providing a broader context for downstream reasoning and decision-making.
@@
SMILES Generation. Constructing a diverse and chemically relevant seed molecule pool is essential for effective early-stage virtual screening. AgentD generates seed molecules in SMILES format using external generative models. In this study, we incorporate two such models: REINVENT 36, 37, which supports de novo molecule generation without requiring input SMILES, and Mol2Mol 38, which performs conditional generation to produce molecules structurally similar to a given input SMILES. This dual capability enables both exploration and exploitation in the molecular search space.
Property Prediction. For each candidate molecule, AgentD predicts key pharmacologically relevant properties, including ADMET (absorption, distribution, metabolism, excretion, and toxicity) profiles and binding affinity (e.g., pKd). ADMET prediction is performed using the Deep-PK API 39, which accepts SMILES strings as input. For binding affinity estimation, the BAPULM model 40 is used, which operates on both the SMILES and the protein’s FASTA sequence. These predictions help prioritize compounds based on both efficacy and safety.
Molecule Refinement. Based on the predicted properties, AgentD can identify molecular shortcomings such as toxicity or poor permeability, and propose targeted structural modifications to improve attributes like solubility and metabolic stability. This SMILES refinement is carried out solely through the LLM’s internal reasoning and built-in chemical knowledge, without additional model-based tools 33.
Structure Generation. AgentD can generate 3D structures of protein–ligand complexes using Boltz-2 as an external tool 41, 42. This process produces candidate complex structures along with associated binding metrics such as IC50 values and inhibitor probability. These structures can serve as inputs for downstream computational tasks such as docking simulations or molecular dynamics, offering deeper insight into the biophysical interactions of the drug candidates.
@@
Question Answering RAG
@@
*Adapted from the dZiner framework 33.
@@ 3.2 Domain-Specific Question Answering
We evaluated this capability using a representative study by Weller et al., which describes how venetoclax activates the integrated stress response (ISR), leading to NOXA upregulation and MCL-1 inhibition 43. Based on this paper, we formulated domain-specific questions and compared the responses generated by AgentD with RAG to those from the standard GPT-4o model (Figure 4).
AgentD consistently generates more detailed and mechanistically accurate responses. For example, when asked about the synergy mechanism between venetoclax and proteasome inhibitors, the RAG-augmented answer includes key elements such as ISR activation, ATF3/ATF4-mediated transcription of NOXA, and downstream effects involving BIM, BAX/BAK activation, and mitochondrial membrane permeabilization—closely aligning with the mechanistic explanation presented in the source paper (Figure 4a). In contrast, GPT-4o provides a more generic explanation, omitting critical components such as the ISR pathway and transcriptional regulators.
@@ 3.3 Seed Molecule Generation
The initial library of molecules serves as a critical foundation for exploring chemical space and identifying candidates for further optimization. To construct this seed molecule pool, AgentD leverages two complementary generation strategies. REINVENT enables de novo molecule generation, allowing for broad and unbiased exploration of chemical space without the need for an input structure. In contrast, Mol2Mol performs conditional generation, producing analogs that are structurally similar to a specified input molecule. In our workflow, the existing drug identified during the data extraction phase—such as venetoclax—is used as input for Mol2Mol, enabling the agent to focus on chemically relevant and biologically meaningful regions of the search space.
@@
Figure 5: t-SNE visualization of chemical space, based on Mordred-derived molecular descriptors 44, including seed molecules generated by AgentD (Mol2Mol, REINVENT), refined molecules after property-aware optimization, 5,000 randomly sampled compounds from ChEMBL, and the reference drug venetoclax.
@@ 3.4 Property-Aware Molecular Refinement
To enhance the quality of the seed molecule pool, AgentD performs property-driven refinement of SMILES structures by identifying and addressing limitations that compromise their drug-likeness. Using the Deep-PK model API 39, the agent predicts 67 ADMET properties including absorption (e.g., Caco-2 $\log P_{\text{app}}$), distribution (e.g., blood–brain barrier permeability), metabolism (e.g., CYP1A2 inhibition), excretion (e.g., drug half-life), and toxicity (e.g., biodegradability, general toxicity, and logD), as well as 7 general physicochemical properties such as lipophilicity and solubility. These properties are critical for evaluating pharmacokinetic profiles and safety. A complete list of predicted ADMET properties is provided in Table S1. In parallel, AgentD uses the BAPULM model 40 to estimate binding affinity (pKd) from the ligand SMILES and the protein FASTA sequence. Details of both models are described in the Methods section.
Based on these predictions, AgentD identifies unfavorable molecular properties relevant to drug development such as low permeability or high toxicity, and proposes targeted structural edits to improve them. A comprehensive list of risk-associated properties is provided in Table S2. We demonstrate this refinement process over two iterations, beginning with an initial set of 100 molecules. These iterations produced 99 and 95 additional valid SMILES, respectively, which were added to the candidate pool. To illustrate how the refinement works in practice, Figure 6a shows a molecule initially flagged for poor permeability due to a low predicted $\log P_{\text{app}}$ value. In response, AgentD suggested replacing a hydroxyl group with a methyl group in the first round, followed by substituting a sulfonamide with an amine in the second - both changes contributed to improved predicted permeability.
@@
Despite these local failures, the overall impact of refinement is positive across the molecule pool. This is supported by increased QED (Quantitative Estimate of Drug-likeness) scores, which reflect overall drug-likeness based on empirical physicochemical properties 45. As shown in Figure 7c, the distribution of high-QED molecules expanded with each refinement iteration: the number of molecules with QED $>$ 0.6 increased from 34 in the original set to 49 after the first update, and to 55 after the second. In addition, more molecules satisfied empirical drug-likeness criteria - including Lipinski’s Rule of Five, Veber’s rule, Ghose’s rule, the Rule of Three, and Oprea’s lead-likeness filter 46, 47, 48, 49, 50. Figure 7d shows a steady rise in the number of molecules passing four or more of these rules, increasing from 29 to 44 and then to 52, indicating improved developability profiles after refinement.
@@
After refining the seed molecule pool through property-aware modifications, we apply empirical filtering criteria to identify high-potential candidates for 3D structural evaluation. Selection is guided by widely accepted drug-likeness rules, including Lipinski’s Rule of Five, Veber’s rule, Ghose’s rule, the Rule of Three, and Oprea’s lead-likeness filter, as detailed in the Table S3 46, 47, 48, 49, 50. It is worth noting that this filtering strategy is illustrative, intended to showcase the agent’s generation capabilities. In our workflow, molecules are shortlisted if they satisfy at least three of these five rules, have a QED score above 0.55, and a predicted pKd value greater than 6.0. It is important to note that this criterion is illustrative and intended to demonstrate the pipeline’s capability for structure generation, and may differ from criteria used in specific therapeutic applications. The molecule shown in Figure 8 meets four of the five drug-likeness rules, has a QED score of 0.68, and a predicted pKd of 6.18, making it a candidate for further analysis. Additional examples are provided in Figure S1.
Once candidates are selected, AgentD uses Boltz-2 to generate 3D protein–ligand complex structures using the ligand’s SMILES and the target protein’s FASTA sequence. These structures provide a foundation for more rigorous downstream evaluations such as molecular docking, MD simulations, and free energy perturbation analyses, to assess binding stability, conformational flexibility, and interaction specificity under biologically relevant conditions. In addition to structural outputs, Boltz-2 returns estimated binding metrics, including IC50 values and inhibitor probabilities. In the example shown, the predicted IC50 and inhibitor probability indicate moderate (not strong) binding affinity, suggesting that further molecular optimization or experimental validation may be necessary.
Figure 8: Protein–ligand structure generation and evaluation. a, A representative 3D complex predicted by Boltz-2, with estimated IC50 and inhibitor probability. b, Radar plot showing empirical rule-based evaluation. Green dotted lines indicate the thresholds for each rule, while the purple solid line represents the properties of the selected molecule. See Methods and Table S3 for property definitions.

### v2

@@ Abstract
We present a modular framework powered by large language models (LLMs) that automates and streamlines key tasks across the early-stage computational drug discovery pipeline. By combining LLM reasoning with domain-specific tools, the framework performs biomedical data retrieval, domain-specific question answering, molecular generation, property prediction, property-aware molecular refinement, and 3D protein–ligand structure generation. In a case study targeting BCL-2 in lymphocytic leukemia, the agent autonomously retrieved relevant biomolecular information, including FASTA sequences, SMILES representations, and literature, and answered mechanistic questions with improved contextual accuracy compared to standard LLMs. It then generated chemically diverse seed molecules and predicted 67 ADMET-related properties, which guided iterative molecular refinement. Across two refinement rounds, the number of molecules with QED >> 0.6 increased from 34 to 55. The number of molecules satisfying empirical drug-likeness filters also rose; for example, compliance with the Ghose filter increased from 32 to 55 within a pool of 100 molecules. The framework also employed Boltz-2 to generate 3D protein–ligand complexes and provide rapid binding affinity estimates for candidate compounds. These results demonstrate that the approach effectively supports molecular screening, prioritization, and structure evaluation. Its modular design enables flexible integration of evolving tools and models, providing a scalable foundation for AI-assisted therapeutic discovery.
@@ 1 Introduction
The discovery of drug-like molecules for treating specific diseases is fundamental to therapeutic innovation and the advancement of public health. However, identifying viable candidates is a highly complex and resource-intensive process. It requires navigating an enormous chemical space, integrating heterogeneous biological and chemical data sources, and conducting iterative experimental validation. Empirically, drug development spans 10-15 years and can cost \$1.5 - 2.5 billion per approved therapy. Much of this burden arises from attrition in Phase II and III clinical trials, where insufficient efficacy or safety issues often lead to late-stage failures 1, 2, 3. A key strategy to reduce this attrition is to identify high-quality candidates as early as possible, thereby improving the likelihood of clinical success. At the same time, the chemical search space is astronomically large, with estimates of more than $10^{60}$ drug-like molecules, making it impossible to explore effectively with traditional trial-and-error approaches 4, 5.
@@
To mitigate these challenges, computational approaches have become essential in early-stage drug development. Methods such as molecular docking 6, 7, 8, quantitative structure–activity relationship (QSAR) modeling 9, 10, 11, and molecular dynamics simulations 12, 13 have significantly improved the speed and accuracy of compound screening and optimization. More recently, machine learning (ML) techniques have expanded this toolkit by enabling predictive modeling of pharmacokinetic and pharmacodynamic properties, de novo molecular generation, and multitask learning across diverse biomedical tasks14, 15, 16, 17. Notably, AlphaFold has achieved high-accuracy protein structure prediction 18, 19, 20, while Graph Neural Networks (GNNs) and transformer-based models improve property prediction 21, 22, 23, 24. In parallel, generative models such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and diffusion models (e.g., JT-VAE, MolGAN, MolDiff) enable de novo or target-specific molecular design 25, 26, 27. Collectively, these methods increase the scalability and automation potential of modern drug discovery workflows.
@@
Despite this progress, most computational methods remain task-specific and require manual orchestration by domain experts. However, drug discovery is inherently a multi-step, interdependent process that requires seamless integration across diverse tasks 28, 29, 30. This fragmented implementation limits the scalability and efficiency of current pipelines. As the demand for faster and more cost-effective drug screening grows, there is an urgent need for unified, intelligent platforms capable of autonomously coordinating these tasks while supporting expert decision-making.
@@
Large Language Models (LLMs) offer a compelling solution to this need. Trained on massive corpora of natural language data, LLMs exhibit strong reasoning capabilities and domain-agnostic knowledge. When augmented with external tools, such as domain-specific plugins, APIs, and software libraries, LLM-based agents can overcome the limitations of general-purpose language models and act as interpretable, flexible controllers of scientific workflows 31. LLM agents have recently been applied to automate diverse aspects of scientific discovery, including experimental design, material synthesis planning, and data analysis 32, 33, 34, 35, 36. For instance, Boiko et al. introduced Coscientist, an LLM agent capable of autonomously planning and executing chemistry experiments, significantly enhancing lab productivity while reducing human intervention 37. Similarly, the dZiner framework leverages LLM agents for molecular design through iterative reasoning and structure-based optimization38.
Furthermore, CACTUS39 demonstrated the use of an LLM to wrap cheminformatics utilities for evaluating properties such as Mwt, LogP, TPSA, and drug-likeness filters, providing a practical but narrowly scoped toolset. In contrast, SciToolAgent40 introduced a knowledge-graph–driven orchestration layer that allows an LLM to select from hundreds of scientific tools across domains ranging from biological to materials, illustrating scalability but lacking a dedicated workflow for therapeutic design. Building on these directions, DrugPilot41 emphasizes parameterized reasoning for multi-stage discovery tasks, though its scope remains tied to a fixed set of model zoo tasks and benchmark datasets, without iterative molecular refinement, limiting its ability to improve candidate quality over successive rounds.
@@ 2.1 Task Modules
Data Extraction. AgentD is capable of retrieving biomedical information from both structured and unstructured sources. Given a query - such as identifying known drugs associated with a specific protein–disease context - the agent retrieves the protein’s FASTA sequence from the UniProt database 42, searches for relevant drug names via the web, and extracts corresponding SMILES representations from the ChEMBL database 43. Additionally, AgentD autonomously constructs keyword-based queries to download relevant open-access scientific literature from Semantic Scholar, providing a broader context for downstream reasoning and decision-making.
@@
SMILES Generation. Constructing a diverse and chemically relevant seed molecule pool is essential for effective early-stage virtual screening. AgentD generates seed molecules in SMILES format using external generative models. In this study, we incorporate two such models: REINVENT 44, 45, which supports de novo molecule generation without requiring input SMILES, and Mol2Mol 46, which performs conditional generation to produce molecules structurally similar to a given input SMILES. This dual capability enables both exploration and exploitation in the molecular search space.
Property Prediction. For each candidate molecule, AgentD predicts key pharmacologically relevant properties, including ADMET (absorption, distribution, metabolism, excretion, and toxicity) profiles and binding affinity (e.g., $pK_{d}$). ADMET prediction is performed using the Deep-PK API 47, which accepts SMILES strings as input. For binding affinity estimation, the BAPULM model 48 is used, which operates on both the SMILES and the protein’s FASTA sequence. These predictions help prioritize compounds based on both efficacy and safety.
Molecule Refinement. Based on the predicted properties, AgentD can identify molecular shortcomings such as toxicity or poor permeability, and propose targeted structural modifications to improve attributes like solubility and metabolic stability. This SMILES refinement is carried out solely through the LLM’s internal reasoning and built-in chemical knowledge, without additional model-based tools 38.
Structure Generation. AgentD can generate 3D structures of protein–ligand complexes using Boltz-2 as an external tool 49, 50. This process produces candidate complex structures along with associated binding metrics such as IC50 values and inhibitor probability. These structures can serve as inputs for downstream computational tasks such as docking simulations or molecular dynamics, offering deeper insight into the biophysical interactions of the drug candidates.
@@
Question Answering RAG, LLM internal knowledge and reasoning
@@
*Adapted from the dZiner framework 38.
@@ 3.2 Domain-Specific Question Answering
We evaluated this capability using a representative study by Weller et al., which describes how venetoclax activates the integrated stress response (ISR), leading to NOXA upregulation and MCL-1 inhibition 51. Based on this paper, we formulated domain-specific questions and compared the responses generated by AgentD with RAG to those from the standard GPT-4o model (Figure 4).
AgentD consistently generates more detailed and mechanistically accurate responses. For example, when asked about the synergy mechanism between venetoclax and proteasome inhibitors, the RAG-augmented answer includes key elements such as ISR activation, ATF3/ATF4-mediated transcription of NOXA, and downstream effects involving BIM, BAX/BAK activation, and mitochondrial membrane permeabilization,closely aligning with the mechanistic explanation presented in the source paper (Figure 4a). In contrast, GPT-4o provides a more generic explanation, omitting critical components such as the ISR pathway and transcriptional regulators.
@@
Table 2: Comparison of BERTScores between AgentD (with RAG) and standard GPT-4o (without RAG). Values are averaged across six question–answer pairs. Results for two representative pairs are shown in Figure 4, and additional examples are provided in Supplementary Information S1.
Method Precision Recall F1-score
AgentD (w/ RAG) 0.720 $\pm$ 0.048 0.687 $\pm$ 0.057 0.702 $\pm$ 0.041
Standard GPT4o (w/o RAG) 0.602 $\pm$ 0.023 0.669 $\pm$ 0.033 0.633 $\pm$ 0.022
To quantitatively assess these observed differences, we measured the semantic similarity between responses generated with and without RAG and human-annotated reference answers derived from the literature using BERTScore 52. The reference answers were curated by the authors through careful reading of the relevant source papers 51, 53. BERTScore quantifies contextual similarity by comparing the embeddings of model-generated and reference texts. We employed SciBERT 54 as the embedding encoder because it is pre-trained on a large corpus of scientific publications from Semantic Scholar, capturing richer representations of biomedical and technical terminology than general-domain encoders such as RoBERTa 41. This setup provides a biologically meaningful and context-aware framework for evaluating the factual basis of model outputs in literature-based question-answering.
As shown in Table 2, AgentD with RAG achieved notably higher semantic fidelity to the reference literature (F1 = 0.70 $\pm$ 0.04) compared to standard GPT-4o (F1 = 0.63 $\pm$ 0.02), with the largest improvement observed in Precision—indicating that RAG effectively filters unsupported or overly generic content. Specifically, for the question shown in Figure 4a, the BERTScores obtained with RAG were Precision = 0.675, Recall = 0.697, and F1 = 0.686, while the corresponding scores without RAG were 0.616, 0.666, and 0.640, respectively. Similarly, for the question in Figure 4b, the BERTScores with RAG were Precision = 0.616, Recall = 0.666, and F1 = 0.640, compared to 0.623, 0.694, and 0.660 without RAG. Consistent with other test cases, these examples demonstrate that incorporating RAG leads to improved semantic alignment and factual grounding.
@@ 3.3 Seed Molecule Generation
The initial library of molecules serves as a critical foundation for exploring chemical space and identifying candidates for further optimization. To construct this seed molecule pool, AgentD leverages two complementary generation strategies. REINVENT enables de novo molecule generation, allowing for broad and unbiased exploration of chemical space without the need for an input structure. In contrast, Mol2Mol performs conditional generation, producing analogs that are structurally similar to a specified input molecule. In our workflow, the existing drug identified during the data extraction phase, such as venetoclax, is used as input for Mol2Mol, enabling the agent to focus on chemically relevant and biologically meaningful regions of the search space.
@@
Figure 5: t-SNE visualization of chemical space, based on Mordred-derived molecular descriptors 55, including seed molecules generated by AgentD (Mol2Mol, REINVENT), refined molecules after property-aware optimization, 5,000 randomly sampled compounds from ChEMBL, and the reference drug venetoclax.
@@ 3.4 Property-Aware Molecular Refinement
To enhance the quality of the seed molecule pool, AgentD performs property-driven refinement of SMILES structures by identifying and addressing limitations that compromise their drug-likeness. Using the Deep-PK model API 47, the agent predicts 67 ADMET properties including absorption (e.g., Caco-2 $\log P_{\text{app}}$), distribution (e.g., blood–brain barrier permeability), metabolism (e.g., CYP1A2 inhibition), excretion (e.g., drug half-life), and toxicity (e.g., biodegradability, general toxicity, and logD), as well as 7 general physicochemical properties such as lipophilicity and solubility. These properties are critical for evaluating pharmacokinetic profiles and safety. A complete list of predicted ADMET properties is provided in Table S2. In parallel, AgentD uses the BAPULM model 48 to estimate binding affinity (pKd) from the ligand SMILES and the protein FASTA sequence. Details of both models are described in the Methods section.
Based on these predictions, AgentD identifies unfavorable molecular properties relevant to drug development, such as low permeability or high toxicity, and proposes targeted structural edits to improve them. A comprehensive list of risk-associated properties is provided in Table S3. We demonstrate this refinement process over two iterations, beginning with an initial set of 100 molecules. These iterations produced 99 and 95 additional valid SMILES, respectively, which were added to the candidate pool. To illustrate how the refinement works in practice, Figure 6a shows a molecule initially flagged for poor permeability due to a low predicted $\log P_{\text{app}}$ value. In response, AgentD suggested replacing a hydroxyl group with a methyl group in the first round, followed by substituting a sulfonamide with an amine in the second - both changes contributed to improved predicted permeability.
@@
Despite these local failures, the overall impact of refinement is positive across the molecule pool. This is supported by increased QED (Quantitative Estimate of Drug-likeness) scores, which reflect overall drug-likeness based on empirical physicochemical properties 56. As shown in Figure 7c, the distribution of high-QED molecules expanded with each refinement iteration: the number of molecules with QED >> 0.6 increased from 34 in the original set to 49 after the first update, and to 55 after the second.
As shown in Figure 7d, LLM-based refinement increased the proportion of molecules meeting empirical drug-likeness and lead-likeness criteria, including Lipinski’s Rule of Five, Veber’s rule, Ghose’s filter, and Oprea’s lead-likeness filter 57, 58, 59, 60. These rules assess different aspects of molecular suitability: Lipinski’s and Veber’s rules emphasize oral bioavailability and permeability, Ghose’s filter captures physicochemical ranges observed in marketed drugs, and Oprea’s filter identifies compounds suitable as lead-like starting points for optimization. After refinement, more molecules satisfied Veber’s rule (78 to 86) and Oprea’s filter (60 to 67). The most notable improvement was observed for Ghose’s filter, with the number of passing molecules increasing from 32 to 55 ( 72% increase). These results indicate that refinement helped expand the chemical space toward drug-like and lead-like regions, particularly in terms of Ghose’s statistical descriptors. Compliance with Lipinski’s rule remained largely unchanged; this outcome is further discussed in the Discussion section.
@@
After refining the seed molecule pool through property-aware modifications, we applied empirical filtering criteria to select candidates for 3D structural evaluation. This filtering scheme is intended as a demonstration rather than a definitive selection strategy. Its purpose is not to identify a final drug candidate, but to illustrate the agent’s capability to perform key tasks in the drug discovery pipeline. In our workflow, molecules are shortlisted if they (i) satisfy Oprea’s lead-likeness filter, (ii) pass at least two of the three drug-likeness rules (Lipinski’s Rule of Five, Veber’s rule, and Ghose’s filter), (iii) achieve a QED score above 0.55, and (iv) have a predicted pKd value greater than 6.0.
The molecule shown in Figure 8 satisfies the lead-likeness requirement, passes all three drug-likeness rules, has a QED score of 0.68, and achieves a predicted pKd of 6.18, making it a representative candidate for further analysis. It is important to note that these empirical rules may not always align, and molecules rarely satisfy all of them simultaneously. In fact, only one molecule in our set passed all three rules, which we selected as an illustrative example for 3D visualization. Additional examples are provided in Supplementary Figure S4.
Once candidates are selected, AgentD uses Boltz-2 to generate 3D protein–ligand complex structures using the ligand’s SMILES and the target protein’s FASTA sequence. These structures provide a foundation for more rigorous downstream evaluations, such as molecular docking, MD simulations, and free energy perturbation analyses, to assess binding stability, conformational flexibility, and interaction specificity under biologically relevant conditions. In addition to structural outputs, Boltz-2 returns estimated binding metrics, including logIC50 values and inhibitor probabilities. In the example shown, the predicted IC50 (0.54 µM) and inhibitor probability indicate good sub-micromolar binding affinity, although further molecular optimization or experimental validation may still be necessary.
Figure 8: Protein–ligand structure generation and evaluation. a, A representative 3D complex predicted by Boltz-2, with estimated IC50 and inhibitor probability. b, Radar plot showing empirical rule-based evaluation. Green dotted lines indicate the thresholds for each rule, while the purple solid line represents the properties of the selected molecule. See Methods and Table S4 for criteria.